OpenStudy (anonymous):

please help!! 1 MEDAL AND FAN Which is the equation of the line with slope 3 that contains point (-1, 5)? y – 5 = 3(x + 1) y – 1 = 3(x + 5) y – 5 = 3(x - 1) y – 1 = 3(x – 5)

OpenStudy (isaiah.feynman):

It's y-5=3(x+1). Thus the standard form becomes y=3x+8. Cutie. :)

OpenStudy (anonymous):

THANK VERY MUCH FOR THE ANSWER

OpenStudy (isaiah.feynman):

No problem. :D

OpenStudy (anonymous):

can you please explain your answer

OpenStudy (isaiah.feynman):

I used the slope-point form equation (y-y1=m(x-x1)) I slotted in the point and the slope values and solved for Y.